Circular No. 6400

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

COMET C/1995 O1 (HALE-BOPP) H. Rauer, Observatoire de Paris-Meudon; H. Böhnhardt, University

Observatory, Munich; F. Colas, Bureau des Longitudes; J. Crovisier, Observatoire de Paris-Meudon; and L. Jorda and N. Thomas, Max-Planck-Institut für Aeronomie, report on observations of this comet on Apr. 27 UT with the 1.54-m Danish telescope (+ DFOSC) at the European Southern Observatory: “Spectra (range 350–750 nm) show sublimation of CN(0-0) at 388 nm, C_{2}(0,0) at 516 nm, C_{2}(1,0) at 475 nm, and C_{3} at 405 nm. C_{2} emission extends to at least 7 × 10^{5} km from the nucleus, while CN can be observed up to 2 × 10^{5} km in the much-noisier blue part of the spectrum. A preliminary data reduction for spectra taken on Apr. 27.29 UT gives production rates of log Q(CN) = 26.8 and log Q(C_{2}) = 26.7, using a Haser model with scale length after Cochran et al. (1992, Icarus 98, 151).”

COMET C/1996 J1 (EVANS-DRINKWATER) R. M. West and L. Ramirez, European Southern Observatory, have

obtained a 30-min spectrum (range 370–900 nm, resolution 0.8 nm) of comet C/1996 J1 with the MPI/ESO 2.2-m telescope (+ EFOSC II) at La Silla on May 14.15 UT. The slit (1′′.5 wide, 5′.7 long at p.a. 90^{o}) was centered on the diffuse central condensation, which shows a solar-type spectrum with weak emission lines of CN (388 nm, extending to ∼ 100′′ on either side) as well as C_{2} (516 nm).

Total visual magnitude estimates by A. Hale, Cloudcroft, NM (0.41-m reflector): May 15.20 UT, 13.6; 16.18, 13.5.

EMISSION OBJECT IN CYGNUS P. Garnavich and P. J. Green, Harvard-Smithsonian Center for Astro-

physics, report: “Spectra (range 350–900 nm; resolution 0.8 nm FWHM) of the object reported on IAUC 6387 were obtained with the Multiple-Mirror Telescope on May 12. The spectra (two 900-s exposures) show a featureless continuum rising sharply to the red. The estimated magnitudes are R = 19.7, I = 17.4. Assuming an intrinsic color of B − V = 0.0, the estimated interstellar absorption is 8 magnitudes in the visual band. No emission lines are seen. This implies a flux limit for Hα of < 2 × 10^{−16} erg cm^{−2} s^{−1}. The red continuum, combined with a lack of absorption bands, suggests that this is a heavily obscured early-type star with highly-variable emission-line strength.”

1996 May 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green